OpenStudy (anonymous):

Solve the system y = –x + 6 and y = x – 8 by using graph paper or graphing technology.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y+%3D+%E2%80%93x+%2B+6 add your other equation in to the system to get your second line.

OpenStudy (lalaly):

x - 8 = -x + 6 subtract 6 from -8 x -14 = -x add x to x 2x - 14 = 0 add 14 2x = 14 x = 7 plug in x y = 7 - 8 y= -1 your answer is d.

OpenStudy (anonymous):

(7,-1) (10,2) (2,2) (1.5,-6.5)

OpenStudy (anonymous):

(7,-1)

OpenStudy (anonymous):

thank you